OpenStudy (anonymous):

Hong can staple the programs for graduation in 30 minutes. However, since Eva has an electric stapler, it only takes her 10 minutes. If they work together, how long will it take to staple the programs?

OpenStudy (anonymous):

\[\frac{10\times 30}{10+30}\]

OpenStudy (anonymous):

always works this way. adam can do the job in x minutes, eve can do the job in y minutes there rates are respectively \(\frac{1}{x}\) and \(\frac{1}{y}\) combines rate is \[\frac{1}{x}+\frac{1}{y}=\frac{x+y}{xy}\] and you want to solve \[\frac{x+y}{xy}T=1\] for one job solution is \[T=\frac{xy}{x+y}\]

Directrix (directrix):

@satellite73 --> I've never heard this type problem explained in such an easy to understand way. I hope I can now unlearn my dislike of that type problem because of your explanation. Thanks.

OpenStudy (anonymous):

why thank you!